Find the coordinates of the x-intercept.
2x + y = -6
I'm not sure how to begin this one.....

Answer by nyc_function(2741)

2x + y = -6
2x + 0 = -6
2x = -6
x = -6/2
x = -3
The x-intercept is found at the point (-3,0), where x = -3 and y = 0.